--++++ INAF queue observing, 2018 September 28-29

Observer: R. Carini, A. Gargiulo
Telescope Operator: D. G. Huerta
Support Astronomer: O. Kuhn on call
Instrument: LUCI

Summary

At the sunset the sky is clear , some cirrus on the horizon
instruments: MODS2 red grating is not working
LBC-blue has some issues LUCI1 AO doesn't work.

Night duration (12\xB0 el twilight) 2:30- 12:15 UT = 9 h 45m

Observing time: 6 35m (%)

Weather time loss: h m (0%)

technical time loss 3h 10m (%)

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
dark Antoniucci            
dark Polletta            
dark telluric Cresci B            
telluric V1057 Cygn HD228448 Luci1+ Luci2 ls, Hkspec, zjSpec luci1.0052-0063,luci2.0056-0067  
Antoniucci V1057 Cyg Luci1+ Luci2 ls, Hkspec, zjspec 0.5/0.5 h luci1.0068-0071,luci2.0072-0076, luci1.0072-0078,luci2.0077-0083 completed
telluric Cresci C HIP102190 ls, zjspec      
Cresci C J2310 ls, zjspec 2/2h luci1.0095-009,luci2.0099-0122 luci1.0113-0118,luci2.0117-0123 seeing variable. completed?
telluric V350 Cep HIP111257 Luci1+Luci2 HKspec,zjSpec    
Antoniucci V350Cep Luci1+Luci2 ls, HKspec, zjSpec   luci1.130-0133, luci2.0134-0137 completed
Cresci J2310 ima 0.07/0.07h luci1.0133-0140 0138-0145 completed
Perna RCSGA327 mos, HKSpec 0.54/0.6 h luci2.0178-0190 max airmass 1.6,seeing very good, 0.7

Details

Dark
  • Antoniucci,:
  • dit 40 ndit 1, mer, norm luci1.20180929.0006, luci2.20180929.0010-
  • 15 ndit 1 mer norm
  • dit 2.75 ndit 3 lir int
  • dit 5 ndit 1 lir norm luci1.0019-0023,luci2.0023-0027
  • dit 30 ndit 1 mer norm luci1.0024-0028,luci2.0028-0032
  • dit 20 ndit 1 mer norm luci1,0029-0033,luci2.0033-0037

  • Polletta
  • dit 120 ndit 1 mer norm luci1,0034-0038,luci2.0038-0042
  • dit 5 ndit 6 lir int luci1,0039-0043,luci2.0043-0047

  • Cresci B telluric
  • dit 5 ndit 3 lir int luci1,0044-0048,luci2.0048-0052

02:24 seeing is bad about 2.0'', we start with Antoniucci program
  • telluric HD228448
  • acq luci1.0049-0051,luci2.0053-0055
  • sci luci1.0052-0063,luci2.0056-0067
  • seeing about 1.2''-1.4''

02:46 slew to V1057 Cyg * 02:49 error commit setup

03:08 load again the script
  • acq luci1.0064-0067,luci2.0068-0071
  • the seeing is decreasing
  • sci luci1.0068-0071,luci2.0072-0076
  • 3:22 error in the grating, error committing setup
  • 03:25 continue integration, the ISA fixes the wavelength but the scripts doesn't go on, so we take a spectra manually, skip the point of the script and after that continue.
  • luci1.0072-0078,luci2.0077-0083
  • seeing about 1.0'''

03:38 slew to Cassar\xE1 program
  • telluric HIP 106817
  • 03:41 check pointing
  • 03:47 load the script
  • check pointing again, LUCI1 can not get the guide star, maybe has a problem with the probe. the To try to understand if LUCI1 can be used.

04:57 problem fixed. We slew to Cresci C program, telluric HIP102190
  • 05:03 the LUCI1 and LUCI2 field stop are not aligned, we do it.

05:21 we load again the script, skip the preset but we configure manually the systems
  • acq luci1.0085-0086,luci2.0089-0091
  • sci luci1.0087-0091,luci2.0092-0095
  • seeing 0.65'''

05:30 send te script for the target C (J2310)
  • LUCI1 has problems with the guide star again.

05:40 send again the script
  • preset failed

05:45 load the script again * preset failed again * 05:51 the script goes on * acq luci1.0092-0094,luci2.0096-0098 * sci luci1.0095-009,luci2.0099-0122 * seeing about 0.9''-1.0'' * 07:00 suddenly the seeing increases up to 2.0''' luci1.0113-0118,luci2.0117-0123 seeing bad * 07:10 seeing about 1.2''-1.5'' * luci1.0113-0118,luci2.0117-0122 seeing bad

07:25 slew to Antoniucci program, telluric for V350 Cep, HIP111257
  • send the script
  • 07:28 check pointing
  • 07:36 we send the script again
  • error in the wheel filter
  • failed the guider
  • send again the script
  • 07:40 acq luci1.0119-0121,luci2.0123-0125
  • 07:53 sci==luci1.0122-0125,luci2.0126-0129==

07:59 send the script for the target V350 Cep
  • acq luci1.0126-0129,luci2.0130-0133
  • seeing about 1.4''-1.5''
  • 08:09 sci luci1.0130-0133,luci2.0134-0137
  • seeing variable between 1.2''-1.4'' with jump to 2.0'''
  • 08:18 seeing 1.3''
  • 08:20 error in the grating, we change it manually
  • seeing 1.1''-1.2''

08:28 seeing 1.0 slew to cresci C imaging
  • 08:30 check pointing
  • 08:33 send again
  • luci133-0140, luci2.0138-0145
  • seeing about 1.0''on the image, luci2.0139 and luci.0134 are not good

08:45 slew to Nisini field 3 because the seeing is too variable
  • 08:49error
  • 8:52 send again the script
  • 09:07 problems with Aladin and the calculation of the offset, the offset are huge
  • problem with Aladin, it doesn't load the images, we can not acquire the telluric
  • we try to close and open again the luci
  • we syncronize the RTD panel
  • we find some errors in the scripts but alone they can not explain all the problems.

10:22 send again the telluric, the calculation of the offset is wrong, since was unexspected activated "use fixed Slit Y position"in both LUCI
  • 10:27 Finally we send the offset but now the target is too near the moon.

10:45 seeing is very good, about 0.8'', we slew to Perna target
  • preset failed
  • 10:54 send again the script
  • problem with AGW
  • 10:56 send again
  • 10:57 preset canceled
  • the probe of LUCI1 doesn't move, we decide to use only LUCI2. TO deactivates LUCI1.
  • 11:01 send the script with LUCI2
  • 11:02 preset failed
  • 11:06 send the script again with the different guide star
  • acq luci2.0174-0177
  • sci luci2.0178-0190
  • 11:47 seeing 0.7''

12:00 slew the telluric
  • preset failed
  • acq luci2.0191-0193
  • sci =luci2.0194-0197

12:24 we take also the telluric in the most distant slit from the first
  • acq luci2.0198-0210
  • sci luci2.0202-0203

Calibrazione Perna
  • flat HKspec luci1.0170-0179-,luci2.0204-0213
  • arc HKspec luci1.0179-0180,luci2.0214-0215 Ne
  • arc HKspec luci1.0181-0182,luci2.0216-0217 Ar *arc HKspec luci1.0183-0184,luci2.0218-0219 Xe
  • dark luci1.0185-0189,luci2.0220-0224
  • flat zjSpec luci1.0190-0194,luci2.0225-0229
  • arc zjSpec ==luci1.0195-0196,luci2.0230-0231==Ne
  • arc zjSpec ==luci1.0195-0196,luci2.0232-0233==Ar

Pinna
  • dark luci2.0234-0.238
  • flat FeII luci2.-0239-0243
  • dark luci2.0244-0248
  • flat H2 luci2.-0249-0253
  • dark luci2.0254-0258
  • flat Ks luci2.0259-0263

-- %USERSIG{RobertaCarini - 2018-09-28}%

Comments

 
Topic revision: r9 - 01 Oct 2018, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback